Measurement no Density (kg/m3
)
1 3.54
2 3.49
3 3.58
4 3.50
5 3.55
6 3.52
(i) Find average density.
1
Expert's answer
2020-12-14T12:14:34-0500

By definition, the average density is the sum of all measurements divided by the number of measurements. Thus, obtain:


"\rho_{av} = \dfrac{3.54+3.49+3.58+3.50+3.55+3.52}{6} = 3.53\dfrac{kg}{m^3}"

Answer. "3.53\dfrac{kg}{m^3}".